# Ephemerality meets LiDAR-based Lifelong Mapping

### v1

@@ IV-A Experimental Setup
Multi-session Map Alignment. For datasets, we utilize six sequences in LT-ParkingLot [11] and MulRan [15] (DCC01-03 and KAIST01-03 sequences) since each of them includes multiple sequences of overlapping routes with sufficient time intervals. Each of the first sessions is utilized for constructing the base map and the others are treated as new sessions for the alignment. We utilize SC-LIO-SAM [3, 39] when building a map from each session. We choose metrics such as accuracy (AC), RMSE, and chamfer distance (CD), following previous works [42, 12]. Given two aligned point clouds, AC reports the ratio of inlier pairs within a threshold $\sigma_{\text{AC}}$, while RMSE is their root mean squared distance. CD measures the mean squared distance between points, excluding those beyond a threshold $\sigma_{\text{CD}}$. Both $\sigma_{\text{AC}}$ and $\sigma_{\text{CD}}$ are set to 0.5m. For the baselines, we select ICP-based map-to-map registration [43] and multi-session PGO in LT-mapper [11].
Dynamic Object Removal. Similar to [31, 33, 29], evaluation is held on three different sequences from the SemanticKITTI [44] dataset, adopting the evaluation protocol outlined in [45]. The metrics, including preservation rate (PR) and removal rate (RR) as proposed in [31], are calculated at the point level without downsampling the ground truth map to ensure accurate evaluation. We additionally report F1 scores, the harmonic mean of PR and RR. The baselines include state-of-the-art method such as Removert [30], ERASOR [31], DUFOMap [29], and BeautyMap [33].
Map Update. Due to the absence of the labeled datasets for inter-session changes, we conduct a qualitative evaluation using the MulRan [15] (KAIST01,02) and HeLiPR [16] (KAIST04,05) datasets. The four-year time gap between the two datasets allows us to evaluate both short-term and long-term changes.
@@ IV-B Multi-session Map Alignment
Sequence Method Metrics CD $\downarrow$ AC $\uparrow$ RMSE $\downarrow$ LT-ParkingLot ICP [43] 0.152 0.876 0.076 LT-mapper [11] 0.162 0.882 0.080 ELite (Ours) 0.128 0.929 0.063 DCC ICP [43] 0.407 0.692 0.204 LT-mapper [11] 0.360 0.738 0.182 ELite (Ours) 0.220 0.942 0.111 KAIST ICP [43] 0.435 0.641 0.217 LT-mapper [11] 0.333 0.909 0.169 ELite (Ours) 0.240 0.963 0.120
@@
Figure 6: Qualitative comparison of multi-session map alignment in MulRan [15] DCC sequence. While both methods exhibit globally consistent alignment, our method shows better local consistency.
@@
Figure 7: Qualitative comparison of utilizing $\epsilon_{g}$ as a weight in scan-to-map registration on KAIST sequence. $\epsilon_{g}$ can be utilized as a reliable weighting factor during multi-session map alignment, prioritizing long-term static points for registration.
Table. I reports map alignment results, in which our method consistently outperforms the baselines. Although ICP may exhibit decent performance in small-scale environments like LT-ParkingLot, it struggles to register large-scale maps via single transformation and exhibits the worst results in DCC and KAIST. Fig. 6 illustrates the alignment of two session maps within DCC sequence (01 and 02). LT-mapper [11] seeks the globally optimal alignment with PGO, which can yield globally consistent but locally misaligned results. Our method, in contrast, exhibits both globally and locally consistent alignment owing to its bidirectional registration. Fig. 7 illustrates the effectiveness of $\epsilon_{g}$ in the registration. By attending more weights to the static structures in the map, our alignment module can accurately register sessions with long time intervals or a number of dynamic objects.
@@ IV-C Dynamic Object Removal
Figure 8: Qualitative comparison of dynamic object removal in the SemanticKITTI [44] 01 sequence. Dynamic points in (a) are highlighted in red and used as ground truth labels. (b)-(f) show the cleaned maps generated by different methods, with red points indicating the remaining dynamic objects.
Figure 9: The sample scene from the KAIST sequences of MulRan [15] and HeLiPR [16]. The red box indicates short-term changes, where the ephemerality of transient objects, such as parked cars, gradually increases through updates. The blue box highlights long-term changes, such as a newly constructed building between sequences 02 and 04. The initial ephemerality of the building is slightly high, suggesting its potential to be ephemeral. The value decreases as more updates are incorporated, making it a static structure.
Table. II and Fig. 8 present the performance of the dynamic object removal methods, where 0.2 and 0.5 denote the thresholds $\tau_{l}$ for the $\epsilon_{l}$. Our approach with $\tau_{l}=0.5$ steadily achieves the highest or comparable performance across most of the sequences compared to the baselines. Additionally, results for $\tau_{l}=0.2$ show a significant increase in RR with a minor trade-off in PR. This demonstrates the flexibility of our method, allowing adjustments based on specific requirements—whether to prioritize dynamic object removal or static points preservation.
We also report the results from two pose sources: KITTI [47] poses and SuMa [46] poses from the SemanticKITTI [44] dataset. As shown in Table. II, all methods show a notable performance drop with KITTI poses compared to SuMa poses, highlighting the impact of pose accuracy. Baseline methods, which rely on spatial quantization, suffer performance degradation due to their sensitivity to pose errors. In contrast, our method remains robust to pose errors by avoiding spatial quantization and utilizing an iterative point-wise update scheme.
@@ IV-D Map Update
Figure 10: Qualitative comparison of representations for inter-session changes. While diff map [11] represents changes with a binary approach, our delta map representation can impose more weights on object-wise changes based on the objectness factor $\gamma$ (with darker colors indicating higher values). This representation can effectively mitigate measurement errors, particularly those prevalent around high-frequency surfaces such as trees.
Fig. 9 illustrates the results of our map update module. The top row shows our lifelong maps updated with each sequential session. After multiple updates, the ephemerality of transient objects, such as parked cars (red boxes), increases, reflecting their dynamic nature. Conversely, ephemerality of static objects, such as newly constructed buildings (blue boxes), gradually decreases, demonstrating their permanence. This highlights that our map update module effectively handles static and ephemeral objects through two-stage ephemerality propagation. Additionally, ELite provides a static map based on a user-adjustable threshold $\tau_{g}$. A low $\tau_{g}$ value results in a map that includes only long-term static structures, while a higher $\tau_{g}$ value incorporates some ephemeral objects, such as parked cars from the current session. This threshold serves as a control knob, enabling users to customize the static map to their needs.
Fig. 10 demonstrates the effectiveness of our delta map compared to the diff map [11]. The diff map provides binary information on point appearance or disappearance, treating true changes and measurement errors—such as tree leaves—identically. In contrast, through the consideration of objectness factor $\gamma$, our delta map is able to accurately distinguish meaningful changes from errors and reflect the characteristics of the real world.
IV-E Potential Down-stream Tasks
Using delta maps, we can generate a heatmap representing the frequency of environmental changes. As shown in Fig. 11, ephemeral objects exhibit high values, effectively highlighting regions with frequent changes. This heatmap can assist robot navigation by identifying regions with a high likelihood of ephemeral objects. Additionally, with a sufficient number of updates, time-domain analysis similar to [24] can be performed. Such analysis can also be leveraged in updating (7) to reflect more region-specific characteristics, which is planned to be handled in our future works.
Figure 11: Left: Static map constructed from the LT-ParkingLot dataset. Right: Heatmap generated from delta maps after updates from six sessions, overlaid with the left. The higher values on the heatmap indicate the higher likelihood of ephemeral objects. This information can be useful for optimizing robot navigation within the space.

### v2

@@ IV-A Experimental Setup
[…]
Dynamic Object Removal. Following [31, 33, 29], we evaluate three different sequences from the SemanticKITTI [44] dataset, adopting the protocol in [45]. We use Preservation Rate (PR) and Removal Rate (RR) [31] at the point level without downsampling the ground truth map to ensure accuracy. We also report the F1 score, the harmonic mean of PR and RR. Baselines include state-of-the-art methods such as Removert [30], ERASOR [31], DUFOMap [29], and BeautyMap [33].
Map Update. Since no labeled dataset exists for inter-session changes, we conduct a qualitative evaluation using the MulRan [15] (KAIST01, 02) and HeLiPR [16] (KAIST04, 05) datasets. With a four-year gap between these datasets, we can observe both short-term and long-term changes.
@@ IV-B Multi-session Map Alignment
Sequence Method Metrics AC $\uparrow$ RMSE $\downarrow$ CD $\downarrow$ LT-ParkingLot ICP [43] 0.962 0.117 0.194 LT-mapper [11] 0.968 0.121 0.175 ELite (Ours) 0.969 0.090 0.133 DCC ICP [43] 0.692 0.204 0.272 LT-mapper [11] 0.738 0.182 0.306 ELite (Ours) 0.942 0.111 0.162 KAIST ICP [43] 0.641 0.218 0.256 LT-mapper [11] 0.909 0.169 0.315 ELite (Ours) 0.963 0.120 0.184
@@
Figure 6: Qualitative comparison of multi-session map alignment on the DCC sequence in MulRan [15]. Both methods exhibit globally consistent alignment, but our method demonstrates superior local consistency.
@@
Figure 7: Qualitative comparison on the KAIST sequence, showing the advantage of using $\epsilon_{g}$ as a weight for scan-to-map registration. By prioritizing long-term static structures, our method robustly aligns sessions even with significant time gaps or dynamic objects.
Table. I shows our map alignment performance, where our method consistently outperforms the baselines. While baselines can perform reasonably well in small-scale settings such as LT-ParkingLot, it struggles to register large-scale environments such as DCC and KAIST. Fig. 6 compares two session maps in the DCC sequence (01 and 02). LT-mapper [11], which leverages multi-session PGO, achieves globally consistent but locally misaligned results. In contrast, our bidirectional registration yields both global and local consistency. Fig. 7 shows the effectiveness of weighting scan-to-map registration with $\epsilon_{g}$. By assigning higher weights to static structures, our alignment module accurately registers sessions separated by huge time intervals or containing many dynamic objects.
@@ IV-C Dynamic Object Removal
Figure 8: Qualitative comparison of dynamic object removal in the SemanticKITTI [44] 01 sequence. The ground truth dynamic points in (a) are shown in red. Figures (b)-(f) depict the cleaned maps produced by each method, with red points indicating remaining dynamic objects.
Table. II and Fig. 8 illustrate the performance of various dynamic object removal methods. We report results for two thresholds of $\epsilon_{l}$: 0.2 and 0.5. Our approach at $\tau_{l}=0.5$ achieves the highest or comparable performance on most sequences. Lowering the threshold to $\tau_{l}=0.2$ significantly increases RR with only a moderate sacrifice in PR, showing the flexibility to prioritize either aggressive removal of dynamic objects or better preservation of static points.
We also evaluate two different pose estimation sources: KITTI poses [47] and SuMa poses [46] from SemanticKITTI [44]. As noted in Table. II, all methods perform worse when using KITTI poses, indicating the importance of accurate pose estimation. Spatial quantization-based methods degrade further with less accurate poses, while our point-wise iterative update remains comparatively robust.
@@ IV-D Map Update
Figure 9: Sample scene from the KAIST sequences of MulRan [15] and HeLiPR [16]. The red box highlights short-term changes, such as parked cars whose ephemerality increases over time. The blue box highlights long-term changes, including a newly constructed building between 02 and 04. While its initial ephemerality is slightly high, indicating potential transience, it decreases as more updates confirm its permanence.
Fig. 9 shows the results of our map update module. The top row depicts the evolving lifelong map over several sessions. After multiple updates, transient objects like parked cars (red box) exhibit increasing ephemerality, aligning with their dynamic nature. Meanwhile, newly built structures (blue box) see a gradual decrease in ephemerality, reflecting long-term permanence. Thus, two-stage ephemerality propagation effectively distinguishes static from ephemeral objects. ELite can also produce a static map by filtering points whose $\epsilon_{g}$ is below a user-defined threshold $\tau_{g}$. A smaller $\tau_{g}$ yields a map retaining only long-term static structures, whereas a higher $\tau_{g}$ preserves moderately ephemeral objects (e.g., currently parked cars). This threshold acts as a convenient way to tailor the static map for specific needs.
@@
Figure 10: Qualitative comparison of inter-session change representations. The diff map [11] uses a binary appearance/disappearance scheme, treating true changes and measurement noise equally. In contrast, our delta map includes an objectness factor $\gamma$ (darker colors correspond to higher values), which assigns greater weights to object-level changes and suppresses noise.
Fig. 10 compares the diff map [11] to our delta map. While the diff map provides only binary information about point appearance or disappearance, our delta map introduces an objectness factor $\gamma$ to accurately separate meaningful changes from artifacts, thus enabling more robust map updates.
IV-E Potential Downstream Tasks
Using delta maps, we can create a heatmap indicating how frequently each region undergoes change. As depicted in Fig. 11, ephemeral objects tend to yield high heatmap values, identifying areas prone to frequent changes—a potentially valuable insight for optimizing robot navigation. Moreover, with enough updates, time-domain analysis similar to [24] can be applied, and this can be integrated into (7) to reflect more region-specific dynamics. We plan to explore this direction in our future work.
Figure 11: Left: Static map constructed from the LT-ParkingLot dataset. Right: Heatmap of frequently changing points after updates from six sessions, overlaid on the static map. Higher heatmap values indicate a stronger likelihood of ephemeral objects, aiding in navigation or planning.